My problem is a reading problem, it is as stated:
There is a strange store that sells 2 apples for $4, and 5 apples for $7. Assuming that the relationship is linear, how much do you think 10 apples will cost in that store?
I believe that they are asking me to add by three on each side, for example:
2+3=5 apples so 5+3=8 apples so 8+3=11 apples which is 1 more apple than I need. $4+$3=$7 so $7+$3=$10 so $10+$3=$13 which would be for 11 apples instead of 10. I am so confused, I do know that I would not shop at that store! Please help! As you can tell I haven't had a math class in 20 years!

Answer by ptaylor(2198)

N=number of apples
C=cost
N C
2 $4 C-N=2
5 $7 C-N=2
10 ? C-N=2
C=N+2
C=10+2=$12
Cost of 10 apples in this store would be $12
Looks to me like $12
Happy Holidays----ptaylor